# DN-4DGS: Denoised Deformable Network with Temporal-Spatial Aggregation for Dynamic Scene Rendering

### v1

@@ Abstract
Dynamic scenes rendering is an intriguing yet challenging problem. Although current methods based on NeRF have achieved satisfactory performance, they still can not reach real-time levels. Recently, 3D Gaussian Splatting (3DGS) has garnered researchers’ attention due to their outstanding rendering quality and real-time speed. Therefore, a new paradigm has been proposed: defining a canonical 3D gaussians and deforming it to individual frames in deformable fields. However, since the coordinates of canonical 3D gaussians are filled with noise, which can transfer noise into the deformable fields, and there is currently no method that adequately considers the aggregation of 4D information. Therefore, we propose Denoised Deformable Network with Temporal-Spatial Aggregation for Dynamic Scene Rendering (DN-4DGS). Specifically, a Noise Suppression Strategy is introduced to change the distribution of the coordinates of the canonical 3D gaussians and suppress noise. Additionally, a Decoupled Temporal-Spatial Aggregation Module is designed to aggregate information from adjacent points and frames. Extensive experiments on various real-world datasets demonstrate that our method achieves state-of-the-art rendering quality under a real-time level.
Figure 1: (a) The visualization results on PlenopticVideo [1] dataset. (b) The visualization results on HyperNeRF [2] dataset. The numbers below the images represent PSNR.
@@ 1 Introduction
Dynamic scene reconstruction from single or multi-view videos is a crucial task in computer vision, with applications such as VR/AR [3; 4], 3D perception [5; 6; 7], movie production [8], etc. Neural Radiance Fields (NeRF) [9] offer a promising approach by representing scenes with implicit functions derived from multi-view inputs. By incorporating time as an additional input [10; 1], NeRF enables dynamic scene rendering. However, the original NeRF model suffers from significant training and rendering costs, attributed to the high number of points sampled per camera ray and volume rendering.
Figure 2: Comparison of our render visualization with 4DGaussian [11]. The results are rendered on HyperNeRF [2] dataset and use the point cloud provided by HyperNeRF for Gaussian initialization ($Sparse$ $Init$). Image 1: canonical 3D gaussians generated by 4DGaussian. Image 2: deformable 3D gaussians generated by 4DGaussian. Image 3: canonical 3D gaussians generated by our method. Image 4: deformable 3D gaussians after the first stage. Image 5: deformable 3D gaussians after the second stage. Image 6: ground truth. The yellow box emphasizes that through a two-stage deformation process, our method can produce higher-quality rendering results.
Recently, the emerging 3D Gaussian Splatting (3DGS) [12] has significantly increased rendering speed to a real-time level compared to NeRFs by employing a differentiable rasterizer for 3D Gaussian primitives. 3DGS directly optimizes the parameters of 3D gaussians (position, opacity, anisotropic covariance, and spherical harmonics (SH) coefficients) and renders them through projection and $\alpha$-blending. Given the explicit expression nature of 3DGS, recent studies [11; 13; 14] represent dynamic scenes by defining a canonical 3D gaussians and deforming it to individual frames in deformable fields. Specifically, during the execution of the deformation field, the coordinates $xyz$ along with time $t$ are used as input, and the output corresponds to the changes in Gaussian properties. It is noteworthy that the existing methods, when designing deformable networks, either directly map the 4D coordinates of each input point to the latent space using MLPs [13; 14], or use HexPlane [15] to interpolate a series of learnable embeddings to obtain the latent of each point [11].
@@
To address the aforementioned issues, we propose Denoised Deformable Network with Temporal-Spatial Aggregation for Dynamic Scene Rendering (DN-4DGS), primarily consisting of two components: the Noise Suppression Strategy (NSS) and the Decoupled Temporal-Spatial Aggregation Module (DTS). To address the initial issue, the design of NSS incorporates two deformation operations. The first deformation operation is a standard deformation. It takes the coordinates $xyz$ of the canonical 3D gaussians and time $t$ as input and outputs corresponding coordinate deformations $\Delta x,\Delta y,\Delta z$. The second deformation builds upon the first by adding $\Delta x,\Delta y,\Delta z$ to the original $xyz$, creating a modified set of coordinates that is then input into a new feature extraction network. The entire process is illustrated by image 3, 4 and 5 in Figure 2. This strategy achieves a successful alteration of the distribution of coordinates $xyz$ through the initial coordinate deformation, resulting in noise reduction and the generation of a more accurate deformation field. It’s worth noting that during the early stage of training, we only perform the first deformation operation. Only after achieving acceptable results with the first deformation operation do we proceed to the second deformation operation to further enhance accuracy. To address the second problem, we design the DTS. The reason we decouple spatial-temporal aggregation is due to the presence of noise in the coordinates of the canonical 3D gaussians. If we directly perform spatial aggregation on the coordinates of the canonical 3D gaussians, the noise information is inevitably amplified after a series of aggregation operations such as k-nearest neighbors (KNN) [16], significantly affecting the results of the deformation field. Therefore, based on the first design NSS, we conduct spatial aggregation during the second deformation operation. Considering that temporal information is unrelated to the canonical 3D gaussians, temporal aggregation can be directly incorporated into the first deformation operation to enhance feature extraction capabilities. In order to reduce computational overhead and considering that temporal information has already been effectively extracted in the first deformation operation, we do not perform temporal aggregation in the second deformation operation. In conclusion, our main contributions are outlined as follows:
@@ 2 Related Work
Dynamic NeRF. Novel view synthesis has been a hot topic in academia for several years. NeRF [9] models static scenes implicitly using MLPs, and numerous studies [17; 18; 19; 2; 10; 20; 21] have extended its application to dynamic scenes through a canonical 3D grid structure and a deformation field. HyperNeRF [2] models object topology deformation using higher-dimensional inputs, while DyNeRF [1] employs time-conditioned NeRF to represent a 4D scene. However, these approaches, based on vanilla NeRF, suffer from high computational costs due to ray point sampling and volume rendering. To address this issue, several acceleration methods [22; 23; 24; 25; 26; 27; 15; 28; 29; 30; 31; 32] have been proposed for rendering dynamic scenes. DeVRF [22] introduces a grid representation, while IBR-based methods [24; 25] utilize multi-camera information for improved quality and efficiency. TensorRF [33] adopts multiple planes as explicit representations for direct dynamic scene modeling. Recent approaches such as K-Planes [34], Tensor4D [30], and HexPlane [15] have also been proposed. NeRFPlayer [35] introduces a unified streaming representation for both grid-based and plane-based methods, utilizing separate models to differentiate static and dynamic scene components, albeit at the cost of slow rendering times. HyperReel [36] suggests a flexible sampling network coupled with two planes for dynamic scene representation. Despite the improvements in training and rendering speed achieved by these methods, they still fall short of meeting real-time requirements.
Dynamic Gaussian Splatting. Recently, 3D Gaussian Splatting (3DGS) [12] has garnered increasing attention from researchers due to its superior rendering quality and real-time rendering speed. The method employs a soft point representation with attributes including position, rotation, density, and radiance, and utilizes differentiable point-based rendering for scene optimization. Soon after, several concurrent works [11; 13; 14; 37] have adapted 3D Gaussians for dynamic scenes. These methods represent dynamic scenes by establishing a canonical 3DGS and deforming it to individual frames using deformable fields. Yang et al. [13] predict per-Gaussian offsets using an additional MLP on canonical 3D gaussians, while Wu et al. [11] substitute the MLP with multi-resolution HexPlanes [15] and a lightweight MLP. Our work introduces the Noise Suppression Strategy to change the distribution of the coordinates of the canonical 3D gaussians and generate a more precise deformation field. Additionally, for better aggregation of temporal-spatial information, we propose the Decoupled Temporal-Spatial Aggregation Module to consolidate information from adjacent points and frames.
@@ 3 Preliminary: 3D Gaussian Splatting
Given images at multiple known viewpoints and timesteps, 3D Gaussian Splatting (3DGS) [12] optimizes a set of attributes (position, opacity, anisotropic covariance and spherical harmonics) via differentiable rasterization. 3DGS can realize high-fidelity rendering of static 3D scenes in real-time.
@@ 4.1 Overview
Our goal is to reconstruct dynamic 3D scenes from single/multi-view videos. Following previous works [11; 13], we represent the geometry and appearance of the dynamic scene using canonical 3D gaussians and model the motion through the deformation fields. An overview of our method is shown in Figure 3. We first describe the details of the Noise Suppression Strategy (NSS) in Section 4.2. Then in Section 4.3, we present the design of the Decoupled Temporal-Spatial Aggregation Module (DTS). Section 4.4 details our optimization process.
@@ 4.3 Decoupled Temporal-Spatial Aggregation Module
Local feature aggregation is very important for 3D point clouds, which can effectively extract local structure information. PointNet++ [16] introduces the set abstraction layer to aggregate information from spatially adjacent points, which has become a fundamental operation in various point cloud tasks [38; 39; 40; 41; 42]. Therefore, to enhance the accuracy of the deformation fields, an intuitive approach is to perform neighbor aggregation for each gaussian.
@@ 4.3.1 Temporal Aggregation Moudle
For each gaussian $G_{t}(i)$, we first input $x,y,z,t$ into the Feature Encoding. Regarding Feature Encoding, we can utilize the MLPs from D3DGS [13] or the HexPlanes from 4DGaussian [11]. After that, we acquire $F_{t}(i)$ at time $t$. Next is a critical step, where we repeat the above-mentioned process to obtain the features $F_{t-1}(i)$ for time $t-1$ and the features $F_{t+1}(i)$ for time $t+1$. It’s worth noting that here, the "1" represents one timestep. Next, similar to PointNet++ [16], we perform the aggregation operation. As illustrated in Figure 4, we merge $F_{t-1}(i)$, $F_{t}(i)$, and $F_{t+1}(i)$ together to form $F(i)\in\mathbb{R}^{3\times 1\times C_{1}}$ and input into a lightweight MLP for channel change. Then, we perform MaxPooling along the first dimension of $F(i)$ to generate $F_{\text{max}}(i)\in\mathbb{R}^{1\times C_{2}}$. Additionally, we introduce a new attribute $\mathcal{Y}_{i}$, which is a learnable embedding. This attribute can provide information independent of coordinates without interference due to adjacency. We have also observed a similar design in a recent work E-D3DGS [43]. Finally, $F_{\text{max}}(i)$, $F_{t}(i)$ and $\mathcal{Y}_{i}$ are concatenated together to generate deformation:
@@
Figure 5: More rendering images of canonical 3D gaussians. Here, $Sparse$ $Init$ refers to using the point cloud provided by the HyperNeRF [2] dataset ($\mathrm{COLMAP_{SFM}}$ [44]) for Gaussian initialization, while $Dense$ $Init$ denotes generating a denser point cloud via $\mathrm{COLMAP_{MVS}}$ [44]. In fact, $Dense$ $Init$ can produce better rendering quality, but due to the need for regenerating, it consumes more computational resources.
@@ 4.4 Optimization
The parameters to be optimized include the deformable network and the attributes of each 3D gaussian $G(i)$: $\mu_{i}$, $\sigma_{i}$, $\Sigma_{i}$, $h_{i}$ and $\mathcal{Y}_{i}$. Following 4DGaussian [11], we use the reconstruction loss $\mathcal{L}_{1}$ and gird-based TV loss [15; 45; 34; 46] $\mathcal{L}_{tv}$ to supervise the training process. Additionally, we add a D-SSIM term $\mathcal{L}_{ssim}$ to improve structural similarity:
@@ 5.1 Experimental Setup
Dataset and Metrics. PlenopticVideo [1] dataset includes 20 multi-view videos, with each scene consisting of either 300 frames, except for the flame salmon scene, which comprises 1200 frames. These scenes encompass a relatively long duration and various movements, with some featuring multiple objects in motion. We utilized PlenopticVideo dataset to observe the capability to capture dynamic areas. Total six scenes (coffee martini, cook spinach, cut roasted beef, flame salmon, flame steak, sear steak) are utilized to train and render. Rendering resolution is set to 1352 $\times$ 1014 HyperNeRF [2] dataset includes videos using two Pixel 3 phones rigidly mounted on a handheld capture rig. We train and render on four scenes (3D Printer, Banana, Broom, Chicken) at a resolution downsampled by a factor of two to 540 $\times$ 960. NeRF-DS [47] dataset consists of seven captured videos (Sieve, Press, Plate, Cup, As, Bell, Basin) with camera pose estimated using colmap [44]. The dataset involves a variety of rigid and non-rigid deformation of various objects. We train and render on the seven scenes. Rendering resolution is set to 480 $\times$ 270.
We report the quality of rendered images using PSNR, SSIM [48], MS-SSIM and LPIPS [49]. Higher PSNR, SSIM and MS-SSIM values and lower LPIPS values indicate better visual quality. To PlenopticVideo dataset, we report PSNR, SSIM and LPIPS (Alex). To HyperNeRF dataset, we report PSNR, SSIM and MS-SSIM. To NeRF-DS dataset, we report PSNR, MS-SSIM and LPIPS (VGG).
Implementation Details. We train our model on a single RTX3090. The optimizer we utilize is Adam [50]. The learning rate is initially set at 1.6e-4, gradually decreasing exponentially to 1.6e-6 by the end of the training process. The learning rate for the voxel grid is initialized at 1.6e-3 and exponentially decays to 1.6e-5. For hyperparameters, we tune $K,\lambda$ as 16, 0.9 respectively. More details will be shown in the Appendix.
@@
Results on PlenopticVideo. Table 1 reports the results on PlenopticVideo dataset. Refer to the Appendix for per-scene details. Due to the incorporation of the Noise Suppression Strategy, which alters the distribution of canonical 3D gaussians coordinates to suppress noise, as well as the utilization of the Decoupled Temporal-Spatial Aggregation Module for feature aggregation, our approach demonstrates superior reconstruction quality across all metrics compared to the baseline (4DGaussian [11]). In fact, PSNR and LPIPS are currently the state-of-the-art metrics. As the table shows, despite our method involving two stages of deformation, resulting in a slight weakening in training time and FPS, it overall meets the requirements for rapid training and real-time demands. Regarding storage, due to the presence of a new attribute $\mathcal{Y}_{i}$, it may slightly exceed the baseline. To vividly illustrate the differences between our method and others, we visualize the qualitative results in Figure 6. From the regions highlighted in red boxes, it is evident that our method can render higher-quality images.
@@
DyNeRF [1] 29.58 - 0.099 1344 hours 0.01 28
NeRFPlayer [35] 30.69 0.909 0.111 6 hours 0.045 -
HyperReel [36] 31.10 0.921 0.096 9 hours 2.0 360
HexPlane-all* [15] 31.70 0.984 0.075 12 hours 0.2 250
KPlanes [34] 31.63 0.964 - 1.8 hours 0.3 309
4DGS [51] 31.19 0.940 0.051 9.5 hours 19.5 8700
E-D3DGS [43] 31.31 0.945 0.037 2 hours 43.1 35
4DGaussian [11] 31.15 0.940 0.049 40 mins 30 90
@@
Results on HyperNeRF. Table 2 reports the results on HyperNeRF dataset. In this dataset, we present results for both $Sparse$ $Init$ and $Dense$ $Init$. $Sparse$ $Init$ refers to using the point cloud provided by the HyperNeRF dataset ($\mathrm{COLMAP_{SFM}}$ [44]) for Gaussian initialization, while $Dense$ $Init$ denotes generating a denser point cloud via $\mathrm{COLMAP_{MVS}}$ [44]. From Table 2, it can be observed that our method outperforms other Gaussian-based methods under both $Sparse$ $Init$ and $Dense$ $Init$ settings. Moreover, under the $Dense$ $Init$ setting, our method achieves the current state-of-the-art performance. More importantly, we find that 4DGaussian is highly sensitive to the sparsity or density of Gaussian initialization. In contrast, our approach benefits from noise suppression and feature aggregation, resulting in a more pronounced performance improvement under the sparse setting. The qualitative results can be observed from Figure 7. From the regions highlighted in red boxes, our method can render higher-quality images, as further supported by PSNR in gray cells.
@@
Nerfies [19] 22.23 - 0.803 $\sim$ hours <1 -
HyperNeRF DS [2] 22.2 0.598 0.811 32 hours <1 -
TiNeuVox-B [45] 24.30 0.616 0.837 30 mins 1 48

### v2

@@ Abstract
Dynamic scenes rendering is an intriguing yet challenging problem. Although current methods based on NeRF have achieved satisfactory performance, they still can not reach real-time levels. Recently, 3D Gaussian Splatting (3DGS) has garnered researchers’ attention due to their outstanding rendering quality and real-time speed. Therefore, a new paradigm has been proposed: defining a canonical 3D gaussians and deforming it to individual frames in deformable fields. However, since the coordinates of canonical 3D gaussians are filled with noise, which can transfer noise into the deformable fields, and there is currently no method that adequately considers the aggregation of 4D information. Therefore, we propose Denoised Deformable Network with Temporal-Spatial Aggregation for Dynamic Scene Rendering (DN-4DGS). Specifically, a Noise Suppression Strategy is introduced to change the distribution of the coordinates of the canonical 3D gaussians and suppress noise. Additionally, a Decoupled Temporal-Spatial Aggregation Module is designed to aggregate information from adjacent points and frames. Extensive experiments on various real-world datasets demonstrate that our method achieves state-of-the-art rendering quality under a real-time level. Code is available at https://github.com/peoplelu/DN-4DGS.
Figure 1: (a) The visualization results on PlenopticVideo li2022neural dataset. (b) The visualization results on HyperNeRF park2021hypernerf dataset. The numbers below the images represent PSNR.
@@ 1 Introduction
Dynamic scene reconstruction from single or multi-view videos is a crucial task in computer vision, with applications such as VR/AR kim2019immersive ; pages2018affordable , 3D perception lu2024bsnet ; li2024mamba24 ; deng2024diff3detr , movie production hisatomi2009method , etc. Neural Radiance Fields (NeRF) mildenhall2021nerf offer a promising approach by representing scenes with implicit functions derived from multi-view inputs. By incorporating time as an additional input pumarola2021dnerf ; li2022neural , NeRF enables dynamic scene rendering. However, the original NeRF model suffers from significant training and rendering costs, attributed to the high number of points sampled per camera ray and volume rendering.
Figure 2: Comparison of our render visualization with 4DGaussian wu20234d . The results are rendered on HyperNeRF park2021hypernerf dataset and use the point cloud provided by HyperNeRF for Gaussian initialization ($Sparse$ $Init$). Image 1: canonical 3D gaussians generated by 4DGaussian. Image 2: deformable 3D gaussians generated by 4DGaussian. Image 3: canonical 3D gaussians generated by our method. Image 4: deformable 3D gaussians after the first stage. Image 5: deformable 3D gaussians after the second stage. Image 6: ground truth. The yellow box emphasizes that through a two-stage deformation process, our method can produce higher-quality rendering results.
Recently, the emerging 3D Gaussian Splatting (3DGS) 3dgs has significantly increased rendering speed to a real-time level compared to NeRFs by employing a differentiable rasterizer for 3D Gaussian primitives. 3DGS directly optimizes the parameters of 3D gaussians (position, opacity, anisotropic covariance, and spherical harmonics (SH) coefficients) and renders them through projection and $\alpha$-blending. Given the explicit expression nature of 3DGS, recent studies wu20234d ; yang2023deformable ; huang2023sc represent dynamic scenes by defining a canonical 3D gaussians and deforming it to individual frames in deformable fields. Specifically, during the execution of the deformation field, the coordinates $xyz$ along with time $t$ are used as input, and the output corresponds to the changes in Gaussian properties. It is noteworthy that the existing methods, when designing deformable networks, either directly map the 4D coordinates of each input point to the latent space using MLPs yang2023deformable ; huang2023sc , or use HexPlane hexplane to interpolate a series of learnable embeddings to obtain the latent of each point wu20234d .
@@
To address the aforementioned issues, we propose Denoised Deformable Network with Temporal-Spatial Aggregation for Dynamic Scene Rendering (DN-4DGS), primarily consisting of two components: the Noise Suppression Strategy (NSS) and the Decoupled Temporal-Spatial Aggregation Module (DTS). To address the initial issue, the design of NSS incorporates two deformation operations. The first deformation operation is a standard deformation. It takes the coordinates $xyz$ of the canonical 3D gaussians and time $t$ as input and outputs corresponding coordinate deformations $\Delta x,\Delta y,\Delta z$. The second deformation builds upon the first by adding $\Delta x,\Delta y,\Delta z$ to the original $xyz$, creating a modified set of coordinates that is then input into a new feature extraction network. The entire process is illustrated by image 3, 4 and 5 in Figure 2. This strategy achieves a successful alteration of the distribution of coordinates $xyz$ through the initial coordinate deformation, resulting in noise reduction and the generation of a more accurate deformation field. It’s worth noting that during the early stage of training, we only perform the first deformation operation. Only after achieving acceptable results with the first deformation operation do we proceed to the second deformation operation to further enhance accuracy. To address the second problem, we design the DTS. The reason we decouple spatial-temporal aggregation is due to the presence of noise in the coordinates of the canonical 3D gaussians. If we directly perform spatial aggregation on the coordinates of the canonical 3D gaussians, the noise information is inevitably amplified after a series of aggregation operations such as k-nearest neighbors (KNN) qi2017pointnet++ , significantly affecting the results of the deformation field. Therefore, based on the first design NSS, we conduct spatial aggregation during the second deformation operation. Considering that temporal information is unrelated to the canonical 3D gaussians, temporal aggregation can be directly incorporated into the first deformation operation to enhance feature extraction capabilities. In order to reduce computational overhead and considering that temporal information has already been effectively extracted in the first deformation operation, we do not perform temporal aggregation in the second deformation operation. In conclusion, our main contributions are outlined as follows:
@@ 2 Related Work
Dynamic NeRF. Novel view synthesis has been a hot topic in academia for several years. NeRF mildenhall2021nerf models static scenes implicitly using MLPs, and numerous studies guo2023forward ; li2021neural ; park2021nerfies ; park2021hypernerf ; pumarola2021dnerf ; tretschk2021non ; xian2021space have extended its application to dynamic scenes through a canonical 3D grid structure and a deformation field. HyperNeRF park2021hypernerf models object topology deformation using higher-dimensional inputs, while DyNeRF li2022neural employs time-conditioned NeRF to represent a 4D scene. However, these approaches, based on vanilla NeRF, suffer from high computational costs due to ray point sampling and volume rendering. To address this issue, several acceleration methods liu2022devrf ; li2023dynibar ; lin2022efficient ; lin2023im4d ; lombardi2021mixture ; peng2023representing ; hexplane ; fang2022fast ; fridovich2023k ; tensor4d ; wang2023mixedvoxels ; wang2023neural have been proposed for rendering dynamic scenes. DeVRF liu2022devrf introduces a grid representation, while IBR-based methods lin2022efficient ; lin2023im4d utilize multi-camera information for improved quality and efficiency. TensorRF chen2022tensorf adopts multiple planes as explicit representations for direct dynamic scene modeling. Recent approaches such as K-Planes kplanes , Tensor4D tensor4d , and HexPlane hexplane have also been proposed. NeRFPlayer song2023nerfplayer introduces a unified streaming representation for both grid-based and plane-based methods, utilizing separate models to differentiate static and dynamic scene components, albeit at the cost of slow rendering times. HyperReel attal2023hyperreel suggests a flexible sampling network coupled with two planes for dynamic scene representation. Despite the improvements in training and rendering speed achieved by these methods, they still fall short of meeting real-time requirements.
Dynamic Gaussian Splatting. Recently, 3D Gaussian Splatting (3DGS) 3dgs has garnered increasing attention from researchers due to its superior rendering quality and real-time rendering speed. The method employs a soft point representation with attributes including position, rotation, density, and radiance, and utilizes differentiable point-based rendering for scene optimization. Soon after, several concurrent works wu20234d ; yang2023deformable ; huang2023sc ; zhu2024motiongs have adapted 3D Gaussians for dynamic scenes. These methods represent dynamic scenes by establishing a canonical 3DGS and deforming it to individual frames using deformable fields. Yang et al. yang2023deformable predict per-Gaussian offsets using an additional MLP on canonical 3D gaussians, while Wu et al. wu20234d substitute the MLP with multi-resolution HexPlanes hexplane and a lightweight MLP. Our work introduces the Noise Suppression Strategy to change the distribution of the coordinates of the canonical 3D gaussians and generate a more precise deformation field. Additionally, for better aggregation of temporal-spatial information, we propose the Decoupled Temporal-Spatial Aggregation Module to consolidate information from adjacent points and frames.
@@ 3 Preliminary: 3D Gaussian Splatting
Given images at multiple known viewpoints and timesteps, 3D Gaussian Splatting (3DGS) 3dgs optimizes a set of attributes (position, opacity, anisotropic covariance and spherical harmonics) via differentiable rasterization. 3DGS can realize high-fidelity rendering of static 3D scenes in real-time.
@@ 4.1 Overview
Our goal is to reconstruct dynamic 3D scenes from single/multi-view videos. Following previous works wu20234d ; yang2023deformable , we represent the geometry and appearance of the dynamic scene using canonical 3D gaussians and model the motion through the deformation fields. An overview of our method is shown in Figure 3. We first describe the details of the Noise Suppression Strategy (NSS) in Section 4.2. Then in Section 4.3, we present the design of the Decoupled Temporal-Spatial Aggregation Module (DTS). Section 4.4 details our optimization process.
@@ 4.3 Decoupled Temporal-Spatial Aggregation Module
Local feature aggregation is very important for 3D point clouds, which can effectively extract local structure information. PointNet++ qi2017pointnet++ introduces the set abstraction layer to aggregate information from spatially adjacent points, which has become a fundamental operation in various point cloud tasks pang2022masked ; wu2022point ; zhao2023divide ; lu2023query ; deng2024unsupervised . Therefore, to enhance the accuracy of the deformation fields, an intuitive approach is to perform neighbor aggregation for each gaussian.
@@ 4.3.1 Temporal Aggregation Moudle
For each gaussian $G_{t}(i)$, we first input $x,y,z,t$ into the Feature Encoding. Regarding Feature Encoding, we can utilize the MLPs from D3DGS yang2023deformable or the HexPlanes from 4DGaussian wu20234d . After that, we acquire $F_{t}(i)$ at time $t$. Next is a critical step, where we repeat the above-mentioned process to obtain the features $F_{t-1}(i)$ for time $t-1$ and the features $F_{t+1}(i)$ for time $t+1$. It’s worth noting that here, the "1" represents one timestep. Next, similar to PointNet++ qi2017pointnet++ , we perform the aggregation operation. As illustrated in Figure 4, we merge $F_{t-1}(i)$, $F_{t}(i)$, and $F_{t+1}(i)$ together to form $F(i)\in\mathbb{R}^{3\times 1\times C_{1}}$ and input into a lightweight MLP for channel change. Then, we perform MaxPooling along the first dimension of $F(i)$ to generate $F_{\text{max}}(i)\in\mathbb{R}^{1\times C_{2}}$. Additionally, we introduce a new attribute $\mathcal{Y}_{i}$, which is a learnable embedding. This attribute can provide information independent of coordinates without interference due to adjacency. We have also observed a similar design in a recent work E-D3DGS bae2024per . Finally, $F_{\text{max}}(i)$, $F_{t}(i)$ and $\mathcal{Y}_{i}$ are concatenated together to generate deformation:
@@
Figure 5: More rendering images of canonical 3D gaussians. Here, $Sparse$ $Init$ refers to using the point cloud provided by the HyperNeRF park2021hypernerf dataset ($\mathrm{COLMAP_{SFM}}$ schonberger2016structure ) for Gaussian initialization, while $Dense$ $Init$ denotes generating a denser point cloud via $\mathrm{COLMAP_{MVS}}$ schonberger2016structure . In fact, $Dense$ $Init$ can produce better rendering quality, but due to the need for regenerating, it consumes more computational resources.
@@ 4.4 Optimization
The parameters to be optimized include the deformable network and the attributes of each 3D gaussian $G(i)$: $\mu_{i}$, $\sigma_{i}$, $\Sigma_{i}$, $h_{i}$ and $\mathcal{Y}_{i}$. Following 4DGaussian wu20234d , we use the reconstruction loss $\mathcal{L}_{1}$ and gird-based TV loss hexplane ; tineuvox ; kplanes ; dvgo $\mathcal{L}_{tv}$ to supervise the training process. Additionally, we add a D-SSIM term $\mathcal{L}_{ssim}$ to improve structural similarity:
@@ 5.1 Experimental Setup
Dataset and Metrics. PlenopticVideo li2022neural dataset includes 20 multi-view videos, with each scene consisting of either 300 frames, except for the flame salmon scene, which comprises 1200 frames. These scenes encompass a relatively long duration and various movements, with some featuring multiple objects in motion. We utilized PlenopticVideo dataset to observe the capability to capture dynamic areas. Total six scenes (coffee martini, cook spinach, cut roasted beef, flame salmon, flame steak, sear steak) are utilized to train and render. Rendering resolution is set to 1352 $\times$ 1014 HyperNeRF park2021hypernerf dataset includes videos using two Pixel 3 phones rigidly mounted on a handheld capture rig. We train and render on four scenes (3D Printer, Banana, Broom, Chicken) at a resolution downsampled by a factor of two to 540 $\times$ 960. NeRF-DS yan2023nerf dataset consists of seven captured videos (Sieve, Press, Plate, Cup, As, Bell, Basin) with camera pose estimated using colmap schonberger2016structure . The dataset involves a variety of rigid and non-rigid deformation of various objects. We train and render on the seven scenes. Rendering resolution is set to 480 $\times$ 270.
We report the quality of rendered images using PSNR, SSIM ssim , MS-SSIM and LPIPS lpips . Higher PSNR, SSIM and MS-SSIM values and lower LPIPS values indicate better visual quality. To PlenopticVideo dataset, we report PSNR, SSIM and LPIPS (Alex). To HyperNeRF dataset, we report PSNR, SSIM and MS-SSIM. To NeRF-DS dataset, we report PSNR, MS-SSIM and LPIPS (VGG).
Implementation Details. We train our model on a single RTX3090. The optimizer we utilize is Adam diederik2014adam . The learning rate is initially set at 1.6e-4, gradually decreasing exponentially to 1.6e-6 by the end of the training process. The learning rate for the voxel grid is initialized at 1.6e-3 and exponentially decays to 1.6e-5. For hyperparameters, we tune $K,\lambda$ as 16, 0.9 respectively. More details will be shown in the Appendix.
@@
Results on PlenopticVideo. Table 1 reports the results on PlenopticVideo dataset. Refer to the Appendix for per-scene details. Due to the incorporation of the Noise Suppression Strategy, which alters the distribution of canonical 3D gaussians coordinates to suppress noise, as well as the utilization of the Decoupled Temporal-Spatial Aggregation Module for feature aggregation, our approach demonstrates superior reconstruction quality across all metrics compared to the baseline (4DGaussian wu20234d ). In fact, PSNR and LPIPS are currently the state-of-the-art metrics. As the table shows, despite our method involving two stages of deformation, resulting in a slight weakening in training time and FPS, it overall meets the requirements for rapid training and real-time demands. Regarding storage, due to the presence of a new attribute $\mathcal{Y}_{i}$, it may slightly exceed the baseline. To vividly illustrate the differences between our method and others, we visualize the qualitative results in Figure 6. From the regions highlighted in red boxes, it is evident that our method can render higher-quality images.
@@
DyNeRF li2022neural 29.58 - 0.099 1344 hours 0.01 28
NeRFPlayer song2023nerfplayer 30.69 0.909 0.111 6 hours 0.045 -
HyperReel attal2023hyperreel 31.10 0.921 0.096 9 hours 2.0 360
HexPlane-all* hexplane 31.70 0.984 0.075 12 hours 0.2 250
KPlanes kplanes 31.63 0.964 - 1.8 hours 0.3 309
4DGS yang2023real 31.19 0.940 0.051 9.5 hours 19.5 8700
E-D3DGS bae2024per 31.31 0.945 0.037 2 hours 43.1 35
4DGaussian wu20234d 31.15 0.940 0.049 40 mins 30 90
@@
Results on HyperNeRF. Table 2 reports the results on HyperNeRF dataset. In this dataset, we present results for both $Sparse$ $Init$ and $Dense$ $Init$. $Sparse$ $Init$ refers to using the point cloud provided by the HyperNeRF dataset ($\mathrm{COLMAP_{SFM}}$ schonberger2016structure ) for Gaussian initialization, while $Dense$ $Init$ denotes generating a denser point cloud via $\mathrm{COLMAP_{MVS}}$ schonberger2016structure . From Table 2, it can be observed that our method outperforms other Gaussian-based methods under both $Sparse$ $Init$ and $Dense$ $Init$ settings. Moreover, under the $Dense$ $Init$ setting, our method achieves the current state-of-the-art performance. More importantly, we find that 4DGaussian is highly sensitive to the sparsity or density of Gaussian initialization. In contrast, our approach benefits from noise suppression and feature aggregation, resulting in a more pronounced performance improvement under the sparse setting. The qualitative results can be observed from Figure 7. From the regions highlighted in red boxes, our method can render higher-quality images, as further supported by PSNR in gray cells.
@@
Nerfies park2021nerfies 22.23 - 0.803 $\sim$ hours <1 -
HyperNeRF DS park2021hypernerf 22.2 0.598 0.811 32 hours <1 -
TiNeuVox-B tineuvox 24.30 0.616 0.837 30 mins 1 48